# Parameter-Efficient Fine-Tuning for Large Models: A Comprehensive Survey

### v1

@@ I Introduction
We categorized all types of PEFT algorithms in Section III according to their computational flow. In Section III-A, we introduce additive algorithms that either introduce additional weight parameters or modify activations. For algorithms that exclusively require fine-tuning with existing parameters, they fall under the category of selective approaches, and their introduction can be found in Section III-B. In Section III-C, we explore reparameterized PEFT, which constructs a (low- dimensional) reparameterization of original model parameters for training while transforms the weights back to maintain the inference speed. Additionally, there exist algorithms that combine the above techniques, and we have classified these as hybrid approaches, elaborating on them in Section III-D. We also investigate strategies for further reducing computational complexity of different PEFT algorithms, including KV-cache management, pruning, quantization, and memory optimization, in Section IV.
@@ III-B Selective PEFT
Rather than additive PEFT, which increases the model complexity by adding more parameters, selective PEFT fine-tunes a subset of the existing parameters to enhance model performance over downstream tasks, as depicted in Figure 4 (b). Specifically, given a model with parameters $\theta=\{\theta_{1},\theta_{2},...,\theta_{n}\}$ where each $\theta_{i}$ denotes an individual model parameter and $n$ represents the total count of these parameters, the process of selective PEFT is represented by applying a binary mask $M=\{m_{1},m_{2},...,m_{n}\}$ to these parameters. Each $m_{i}$ in $M$ is either 0 or 1, indicating whether the corresponding parameter $\theta_{i}$ is selected (1) or not selected (0) for fine-tuning. The updated parameter set $\theta^{\prime}$ after fine-tuning is given by:
@@ III-D Hybrid PEFT
Figure 9: Taxonomy of Efficient PEFT Design
@@ VI-A4 Distributed Metrics
Algorithm 1 KR Algorithm (simplified version)
1:Inputs: LLM module pool F; Total number of layer L; Selection criteria $\epsilon$. Total finetuning epoch E;
[…]
3:Tune all PEFT modules within F for several iterations to collect statistics.
4:Record the changes on $l_{2}$ norms for each module, define the selection criteria.
[…]
6:for $10\leq e\leq E-$ 1 do
7:     for $l\in L$ do
8:         if $l_{2}$ norm change on $l$-$th$ PEFT module less than $\epsilon$ then
9:              Frozen the PEFT module at layer l.
10:         else
11:              Activate $l$-th PEFT module.
12:     Finetune all active PEFT modules.
13:     Recalculate $\epsilon$.
@@ VI-A5 Multi-PEFT Training
VI-B Case study: DLoRA
Due to the scale of LLM, PEFT operations are usually executed in the public environment (e.g., cloud server). This necessitates the sharing of sensitive user data across public environments, thereby raising potential privacy concerns. To tackle these challenges, a solution called DLoRA is proposed. DLoRA enables scalable PEFT operations to be performed collaboratively between the cloud and user devices. The evaluation results demonstrate that DLoRA can significantly reduce the computation and communication workload over the user devices while achieving superior accuracy and privacy protection.
Figure 11: $l_{2}$-norm variation of selected PEFT modules across training iterations over multiple downstream tasks including (a) Arc-Challenge, (b) Social-QA.
DLoRA relies on an algorithm 1 termed Kill and Revive (KR) to reduce user-end computation and cloud and user communication. DLoRA recorded different PEFT modules’ weight changes in different decoders during the tuning session. Figure 11 shows the sampled PEFT module weights over 4 epochs. We observe that some PEFT modules change greatly while others remain stable. DLoRA proposes that edge-side PEFT modules can be killed when their weight changes are not noticeable and be revived when their weight changes are significant.
The cloud and user ends can conduct computation simultaneously. After receiving the activation for a specific PEFT module from the cloud, the user device will start computing output activation for the PEFT module. In the meantime, the cloud ends can continue computing other blocks. The synchronous barrier between users and the cloud is the activation functions (e.g. ReLU) in each decoder. This is because PEFT module computation normally performs residual block before each activation layer.

### v2

@@ I Introduction
We categorized all types of PEFT algorithms in Section III according to their computational flow. In Section III-A, we introduce additive algorithms that either introduce additional weight parameters or modify activations. For algorithms that exclusively require fine-tuning with existing parameters, they fall under the category of selective approaches, and their introduction can be found in Section III-B. In Section III-C, we explore reparameterized PEFT, which constructs a (low- dimensional) reparameterization of original model parameters for training while transforms the weights back to maintain the inference speed. Additionally, there exist algorithms that combine the above techniques, and we have classified these as hybrid approaches, elaborating on them in Section III-D. We also investigate strategies for further reducing the computational complexity of different PEFT algorithms, including KV-cache management, pruning, quantization, and memory optimization, in Section IV.
@@ III-B Selective PEFT
Rather than additive PEFT, which increases the model complexity by adding more parameters, selective PEFT fine-tunes a subset of the existing parameters to enhance model performance over downstream tasks, as depicted in Figure 4 (b).
Specifically, given a model with parameters $\theta=\{\theta_{1},\theta_{2},...,\theta_{n}\}$ where each $\theta_{i}$ denotes an individual model parameter and $n$ represents the total count of these parameters, the process of selective PEFT is represented by applying a binary mask $M=\{m_{1},m_{2},...,m_{n}\}$ to these parameters. Each $m_{i}$ in $M$ is either 0 or 1, indicating whether the corresponding parameter $\theta_{i}$ is selected (1) or not selected (0) for fine-tuning. The updated parameter set $\theta^{\prime}$ after fine-tuning is given by:
@@ III-D Hybrid PEFT
Figure 9: Taxonomy of Efficient PEFT Design.
@@ VI-A5 Multi-PEFT Training
VI-B Case study: Offsite-Tuning
We already know that fine-tuning LLM for downstream tasks is challenging for two reasons: dual privacy concerns between cloud server and data owner, and issues with computational resources and efficiency. Firstly, the privacy of both parties is at risk: the weights of large models are often proprietary and not made public. Sharing data with model owners for fine-tuning can lead to data privacy concerns while providing model weights to data proprietors could compromise the ownership of proprietary models. Secondly, even if downstream users have access to pre-trained weights, the stringent hardware requirements make transfer learning impractical for most end users.
To resolve these two issues, Offsite-Tuning [235] proposes a privacy-preserving and efficient transfer learning framework that enables foundational models to adapt to downstream tasks without the need to access the complete model weights. The key insight of Offsite-Tuning is the cloud provider sends an adapter and an emulator to the data proprietor. Then, with the assistance of the emulator, the data proprietor fine-tunes the adapter. The fine-tuned adapter is then sent back to the cloud side, which integrates it into the complete model, creating a fine-tuned foundational model for downstream users.
Offsite-Tuning safeguards the privacy of data proprietors since they do not need to share their training data directly. It also protects the foundational model owners, as the complete model weights are not shared, and the emulator provided is lossy, with significantly degraded performance. Compared to existing fine-tuning methods that require access to the full model weights, Offsite-Tuning is more resource-efficient because it allows for fine-tuning through a compressed emulator without needing the complete model.